# Internet of Agents: Fundamentals, Applications, and Challenges

### v1

@@ II-D Emerging IoA Application Scenarios
  • Scenario 3: Cross-Domain Agent Networking in Smart City. At the urban scale, IoA fosters dynamic cross-domain collaboration among heterogeneous agents from diverse stakeholders, such as municipal traffic controllers, public safety UAVs, autonomous vehicles, and emergency responder robots. By forming on-demand task-specific teams secured via mutual identity authentication, agents from different entities can rapidly assemble to conduct real-time surveillance, detect anomalies (e.g., fires or traffic incidents), and coordinate multimodal responses. Upon task completion, resources are released automatically, optimizing network load and ensuring scalable and resilient orchestration of city-wide services.
Summary and Lessons Learned: This subsection first surveys existing key multi-agent frameworks, ranging from goal-driven systems (e.g., AutoGPT and BabyAGI) to modular pipelines (e.g., LangChain) and role-based simulations (e.g., MetaGPT), highlighting advances in inter-agent reasoning, planning, and tool integration. We then sketch a technology roadmap, ranging from GUI-driven CUAs and AI smartphones to MCP and A2A protocols and future IoA protocols. Finally, we illustrate three representative IoA application scenarios: P2P agent communication in smart homes, coordination between on-site and external agents in smart factories, and cross-domain networking in smart cities. Collectively, these developments underscore the need for seamless interoperability, agent-native interfaces, and protocol-driven self-organization as the foundation for future IoA.
@@ III-A1 Capability Evaluation of Agents
  • Self-reported Capability Declaration. In IoA, each newcomer agent provides a self-reported capability profile during registration, detailing its foundational model characteristics, available tools, and functional expertise [7, 9]. For instance, an agent might declare its profile as “{model: “GPT-4o”, tools: [“Web Browser”, “Code Executor”], description: “Specialized in QA tasks”}”. While self-reporting facilitates rapid onboarding, it is inherently prone to inaccuracies or overstatements.
@@ III-A3 Capability Retrieval of Agents
TABLE V: Comparison of Agent Capability Retrieval Strategies
@@
Incorporate agents’ capabilities
@@
Summary and Lessons Learned: Capability discovery between agents is a prerequisite for effective task allocation and collaborative orchestration in IoA, including three key components: capability evaluation, capability notification, and capability retrieval. ① Capability evaluation includes self-reported declarations and system-level verification to construct agent capability profiles. While self-reporting facilitates fast registration, it may lead to inaccurate or exaggerated claims; system verification improves reliability but introduces scalability bottlenecks and potential privacy risks. ② Capability notification maintains real-time system awareness through proactive updates, event-triggered mechanisms, and periodic synchronization. Nevertheless, frequent updates in large-scale IoA ecosystem can cause communication overhead and consistency issues, posing challenges to efficiency and scalability. ③ Capability retrieval via traditional matching, semantic retrieval, knowledge-enhanced methods, and agentic-enhanced strategies can improve the accuracy and flexibility of agent selection. Table V provides a comparative summary of representative techniques and their key characteristics.
@@ III-B1 Interaction Structure Design
  • Debate. This mode promotes argumentative collaboration, where agents engage in structured dialogue to challenge, justify, and improve upon each other’s predictions. Each participant updates its response by considering the viewpoints and counterarguments presented by its peers. Empirical studies have shown that such multi-agent debates often yield more truthful and accurate conclusions compared to single-agent reasoning [64, 40].
@@
TABLE VI: Comparison of existing task orchestration methods.
@@
Task Decomposition Rule-based Decomposition Decomposing tasks based on pre-defined logical rules, schemes, or symbolic structures High interpretability; Controllability Limited adaptability to dynamic environments; High manual effort TDAG [98], HM-RAG [99]
@@
Task Allocation Routing-based Allocation Dynamically matching task characteristics with suitable agents or models via optimized routing mechanisms High efficiency, High scalability Depends on the quality of routing strategy; Require additional optimization RouteLLM [101], Hybrid-LLM [102], RouterEval [103]
@@ III-B2 Task Orchestration
Summary and Lessons Learned: In IoA, well-designed interaction structures and task orchestration are foundational to achieve scalable and efficient multi-agent collaboration. On one hand, interaction structures combine both interaction modes and communication topologies. Interaction modes, such as aggregate, reflect, debate, and tool-use, define how agents share, refine, or challenge each other’s outputs, while communication topologies, including chain, star, tree, and graph, shape efficiency, scalability, and coordination dynamics. Recent advances have moved beyond fixed designs, employing learning-based and task-driven approaches that adapt communication patterns and agent roles to contextual needs. On the other hand, task orchestration includes decomposing complex tasks and assigning subtasks to suitable agents. Rule-based and learning-based decomposition techniques offer complementary trade-offs between interpretability and adaptability, while routing-based and self-organizing task allocation strategies balance centralized efficiency and decentralized flexibility. Despite their diversity, these methods share a common goal: to enable dynamic, autonomous, and robust cooperation among heterogeneous agents in complex environments. Table VI summarizes representative orchestration approaches, highlighting their strategies, advantages, and limitations.
TABLE VII: Comparison of Representative Agent Communication Protocols
@@
Large model’s access
@@
Figure 9: The workflow of Anthropic’s MCP.
Figure 10: The workflow of Google’s A2A.
@@ III-C Communication Protocols for IoA
Traditionally, most agent frameworks only consider agents operating within their own ecosystems [7, 52, 41, 53, 37, 54, 47]. Such isolated architecture hinders interoperability, restricts the integration of third-party agents, and fails to reflect real-world scenarios where agents are distributed across heterogeneous devices and contexts. To overcome these fragmentation challenges, recent efforts focus on standardized agent communication protocols to support message exchange, task negotiation, context sharing, identity management, and integration with external tools, forming the backbone of scalable and interoperable IoA ecosystems. Representative agent communication protocols include the model context protocol (MCP) [13], agent-to-agent (A2A) [9], agent network protocol (ANP) [46], AGNTCY [109], and Agora [11], as summarized in Table VII.
III-C1 Anthropic’s MCP
@@
III-C2 Google’s A2A
Unlike the MCP, which focuses on connecting individual LLMs with external data sources and tools, Google’s A2A protocol serves as a networking layer to facilitate seamless interoperability among heterogeneous AI agents. It facilitates direct communication and knowledge sharing for task collaboration. A2A incorporates enterprise-grade authentication and authorization mechanisms compliant with OpenAPI standards by default [9], such as OAuth 2.0 and OpenID Connect.
A2A facilitates communication between a client agent and a remote agent [9], as shown in Fig. 10. The client agent formulates and transmits tasks, while the remote agent responds by executing the tasks or providing relevant information. Agent collaboration begins when the client agent discovers a remote agent by retrieving its agent card, i.e., a standardized JSON file typically hosted at a well-known URL. This agent card contains essential metadata, including agent’s capabilities, skills, endpoint address, and authentication requirements, enabling structured agents discovery across diverse systems. Then, the client agent initiates a task by sending a request containing a unique task ID via universally unique identifier (UUID).The remote agent handles the task according to its declared capabilities. For asynchronous tasks, intermediate server-sent events (SSE) are streamed as the task progresses, providing status updates and partial response. In contrast, synchronous tasks return the final result directly in the response. If additional input is needed, the client sends follow-up messages using the same task ID until the task reaches a terminal state (e.g., completed or failed).
@@ III-C3 Other Protocols
  • Mixture-of-Agents (MoA). MoA [112] enhances agent coordination emulating a “collaborative LLM” through a layered architecture, where each layer contains multiple LLM-based agents with the same model architecture that refine the previous layer’s outputs as auxiliary context.
Summary and Lessons Learned: In IoA systems, standardized communication protocols are vital to enable interoperability among heterogeneous agents, laying the foundation for scalable and context-aware collaboration. Notably, MCP provides agents with structured and standardized access to external tools and resources through unified interfaces, while A2A emphasizes P2P interactions and dynamic agent discovery. In parallel, ANP focuses on decentralized networking; AGNTCY adopts a hybrid architecture to promote standardized inter-agent collaboration; Agora explores agent integration with a balance of efficiency, versatility, and portability; and MoA targets on task-oriented layered agent orchestration. Despite differences in architecture and functionality, these protocols share a common goal: to build thriving, extensible, and trustworthy IoA ecosystems. Table VII provides a comparative summary of their core goals, architectural designs, authentication mechanisms, agent discovery strategies, endorsing entities, advantages, and limitations.
@@ III-D Consensus & Conflict Resolution in IoA
2) How to Align Diverse Agent Reasoning? To overcome the inherent limitations of LLM agents, such as diverse reasoning paths and hallucination, reasoning alignment techniques extend beyond simple majority voting [61] to achieve more reliable consensus.
@@
  • Collective Reasoning: The other promising direction is to utilize the collective power of multiple LLM agents, enabling mutual verification and augmentation. Such collective process enhances inference depth and robustness, particularly for complex, multi-hop reasoning tasks. For instance, He et al. [114] demonstrates that aggregating reasoning from diverse models can improve factual accuracy and approximate correctness, even without ground truth.
3) How to Scale Consensus in Large-Scale IoA? To mitigate scalability issues in consensus formation, hierarchical consensus mechanisms could be employed. Agents are partitioned into local clusters based on network topology or task similarity, with each cluster forms local consensus independently. Then, cluster representatives propagate aggregated results to higher-level coordinators, enabling efficient, scalable, and coherent system-wide decision-making. For instance, Feng et al. [115] propose hierarchical consensus framework that forms global consensus from agents’ local observations via contrastive learning. By incorporating multi-layer consensus and an adaptive attention mechanism, the framework enhances decentralized cooperation and improves scalability.
@@ III-E1 Pricing Models
  • Contribution-aware Pricing rewards agents based on their marginal impact on task outcomes. A common approach employs the Shapley value to quantify each agent’s marginal contribution. Specifically, agents with greater influence over critical task components receive proportionally larger shares of the collective reward, enabling fair profit distribution in multi-agent settings. For instance, Heuillet et al. [59] apply Monte Carlo sampling to estimate Shapley values in a MARL setting to infer and reward cooperative behavior in complex environments.
@@ III-E2 Economic Mechanism Design
TABLE VIII: Comparison of Incentive Mechanism Designs in IoA Systems
@@
a) Incentive Mechanism Design. Typically, agents are rewarded through monetary rewards, e.g., tradable tokens redeemable for computing resources, based on task performance [119]. Prior works on incentives in MAS provides valuable lessons for IoA, including auction theory [120], contract theory [121], game theory [116, 117], and reputation-based mechanisms [122].
@@
  • Auction Theory. In auction mechanisms, agents bid for tasks based on resource availability and task valuation, enabling fair competition and efficient allocation. For instance, You et al. [121] design a stochastic auction mechanism for efficient and privacy-preserving resource allocation among multiple agents.
@@
  • Contract Theory. By specifying performance-based contract items for agents with diverse types, contract-theoretic approaches provide a structured approach to mitigate information asymmetry and incentivize agents to align their actions with the objective of the principal (e.g., task coordinator agent). For instance, Ye et al. [121] employ contract theory and generative diffusion models to optimize multi-dimensional resource allocation under asymmetric information, offering insights for incentive design in IoA systems.
@@
  • Game Theory. Game-theoretic approaches in IoA are divided into non-cooperative and cooperative models games. ① Non-cooperative games model agents as rational and selfish entities whose strategic interactions converge to a Nash equilibrium, at which no agent can unilaterally improve its payoff by deviating. For instance, Geng et al. [116] formulate a non-cooperative game among distributed quadrotor UAVs and derive a generalized Nash equilibrium of the game under communication and physical constraints. ② Cooperative games encourage agents to independently form coalitions, with coalitional benefits distributed fairly among its members via mechanisms such as Shapley values. For instance, Wang et al. [117] model social cluster formation as a cooperative game, where agents collaboratively optimize group strategies to enhance overall utility while ensuring equitable coalition benefit allocation.
@@
  • Reputation mechanisms track agents’ historical performance, prioritizing high-reputation agents for critical tasks and discouraging malicious behavior. For instance, Xu et al. [122] propose a hierarchical reputation framework that selects reliable agents based on reputation scores for secure collaboration in dynamic and uncertain environments.
@@
  • Game-theoretic Penalties. Leveraging principles from game theory, several incentive-compatible penalty schemes have developed to align individual agent behavior with system-wide objectives. For instance, Xu et al. [123] design game-theoretical reward and penalty policies based on agents’ reported beliefs and past performance, thereby improving task quality while reducing latency.
@@
Summary and Lessons Learned: Economic models in IoA align individual agent incentives with collective objectives to sustain long-term collaboration. In IoA pricing models, capability-based pricing charges agents according to resource usage (e.g., token usage), whereas contribution-aware pricing allocates rewards agents proportionally to their marginal contributions. However, accurately estimating contributions in large-scale and dynamic IoA environments is often computationally prohibitive. Beyond pricing, mechanism design integrates incentives and penalties to foster cooperation. Incentive mechanisms grounded in auction, contract, and game theories promote active participation, while reputation mechanisms support long-term collaboration. Table VIII summarizes representative incentive works on agents along with their strengths and limitations. Conversely, penalty mechanisms deter misbehaving agents by imposing costs (e.g., token forfeiture or reputation loss). A key challenge is developing interoperable and fair economic frameworks across heterogeneous IoA platforms with varying protocols and value models.

### v2

@@ II-D Emerging IoA Application Scenarios
  • Scenario 3: Coordination between Agents in an IoA Subnet and and External Agents in Smart Factory. Within a smart factory, the IoA subnet connects on-site manufacturing agents (e.g., robotic arms, autonomous guided vehicles, and digital twin controllers) with external ecosystem participants (such as component suppliers, logistics UAVs, and cloud analytic agents). For instance, an online production supervisor agent can request spare parts through a supplier's digital assistant, assign an autonomous forklift to negotiate pickup timing with a delivery UAV, and authenticate identities via QR-code or device-to-device (D2D) communications. Throughout this process, the network provides digital identity verification and sensing-based path-planning services, ensuring that materials arrive just-in-time and production workflows adapt responsively to supply-chain dynamics.
  • Scenario 4: Energy Management via IoA Subnet and External Agents in Smart Grid. Within smart grid ecosystems, an IoA subnet interconnects domain-specific energy agents, such as renewable generator agents, storage system agents, smart meter agents, and demand controller agents. These agents coordinate locally to balance electricity production, storage, and consumption in real time. At the same time, the subnet interacts with external entities including regional grid operators, energy trading platforms, and forecasting services. For instance, renewable generator agents negotiate surplus electricity allocation with external market agents, while storage agents synchronize with operator agents to provide ancillary services such as peak shaving and frequency regulation. Meanwhile, smart meter and demand controller agents exchange data with forecasting services to anticipate consumption fluctuations and dynamically adjust appliance usage. Through this integration, the IoA subnet ensures resilient, demand-responsive energy distribution, while external collaborations enhance system-wide efficiency and economic optimization.
  • Scenario 5: Cross-Domain Agent Networking in Smart City. At the urban scale, IoA fosters dynamic cross-domain collaboration among heterogeneous agents from diverse stakeholders, such as municipal traffic controllers, public safety UAVs, autonomous vehicles, and emergency responder robots. By forming on-demand task-specific teams secured via mutual identity authentication, agents from different entities can rapidly assemble to conduct real-time surveillance, detect anomalies (e.g., fires or traffic incidents), and coordinate multimodal responses. Upon task completion, resources are released automatically, optimizing network load and ensuring scalable and resilient orchestration of city-wide services.
Summary and Lessons Learned: This subsection first surveys existing multi-agent frameworks, ranging from goal-driven systems (e.g., AutoGPT and BabyAGI) to modular pipelines (e.g., LangChain) and role-based simulations (e.g., MetaGPT), highlighting advances in inter-agent reasoning, planning, and tool integration. We then sketch a technology roadmap, ranging from GUI-driven CUAs and AI smartphones to MCP and A2A protocols and future IoA protocols. Finally, we illustrate several representative IoA application scenarios: P2P agent communication (in smart homes and smart healthcare), coordination between on-site and external agents (in smart factories and smart grids), and cross-domain networking (in smart cities). Collectively, these developments underscore the need for seamless interoperability, agent-native interfaces, and protocol-driven self-organization as the foundation for future IoA.
@@ III-A1 Capability Evaluation of Agents
  • Self-reported Capability Declaration. In IoA, each newcomer agent provides a self-reported capability profile during registration, detailing its foundational model characteristics, available tools, and functional expertise [7, 9]. For instance, an agent might declare its profile as ``{model: ``GPT-4o'', tools: [``Web Browser'', ``Code Executor''], description: ``Specialized in QA tasks''}''. While self-reporting facilitates rapid onboarding, it is inherently prone to inaccuracies or overstatements.
@@ III-A3 Capability Retrieval of Agents
TABLE VII: Comparison of Agent Capability Retrieval Strategies
@@
Incorporate agents' capabilities
@@
Summary and Lessons Learned: Capability discovery between agents is a prerequisite for effective task allocation and collaborative orchestration in IoA, including three key components: capability evaluation, capability notification, and capability retrieval. ① Capability evaluation includes self-reported declarations and system-level verification to construct agent capability profiles. While self-reporting facilitates fast registration, it may lead to inaccurate or exaggerated claims; system verification improves reliability but introduces scalability bottlenecks and potential privacy risks. ② Capability notification maintains real-time system awareness through proactive updates, event-triggered mechanisms, and periodic synchronization. Nevertheless, frequent updates in large-scale IoA ecosystem can cause communication overhead and consistency issues, posing challenges to efficiency and scalability. ③ Capability retrieval via traditional matching, semantic retrieval, knowledge-enhanced methods, and agentic-enhanced strategies can improve the accuracy and flexibility of agent selection. Table VII provides a comparative summary of representative techniques and their key characteristics.
@@ III-B1 Interaction Structure Design
  • Debate. This mode promotes argumentative collaboration, where agents engage in structured dialogue to challenge, justify, and improve upon each other's predictions. Each participant updates its response by considering the viewpoints and counterarguments presented by its peers. Empirical studies have shown that such multi-agent debates often yield more truthful and accurate conclusions compared to single-agent reasoning [66, 40].
@@
TABLE VIII: Comparison of existing task orchestration methods.
@@
Task
Decomposition
Rule-based Decomposition Decomposing tasks based on pre-defined logical rules, schemes, or symbolic structures High interpretability; Controllability Limited adaptability to dynamic environments; High manual effort TDAG [100], HM-RAG [101]
@@
Task
Allocation
Routing-based Allocation Dynamically matching task characteristics with suitable agents or models via optimized routing mechanisms High efficiency, High scalability Depends on the quality of routing strategy; Require additional optimization RouteLLM [103], Hybrid-LLM [104], RouterEval [105]
@@ III-B2 Task Orchestration
Summary and Lessons Learned: In IoA, well-designed interaction structures and task orchestration are foundational to achieve scalable and efficient multi-agent collaboration. On one hand, interaction structures combine both interaction modes and communication topologies. Interaction modes, such as aggregate, reflect, debate, and tool-use, define how agents share, refine, or challenge each other's outputs, while communication topologies, including chain, star, tree, and graph, shape efficiency, scalability, and coordination dynamics. Recent advances have moved beyond fixed designs, employing learning-based and task-driven approaches that adapt communication patterns and agent roles to contextual needs. On the other hand, task orchestration includes decomposing complex tasks and assigning subtasks to suitable agents. Rule-based and learning-based decomposition techniques offer complementary trade-offs between interpretability and adaptability, while routing-based and self-organizing task allocation strategies balance centralized efficiency and decentralized flexibility. Despite their diversity, these methods share a common goal: to enable dynamic, autonomous, and robust cooperation among heterogeneous agents in complex environments. Table VIII summarizes representative orchestration approaches, highlighting their strategies, advantages, and limitations.
TABLE IX: Comparison of Representative Agent Communication Protocols
@@
Large model's access
@@
Figure 9: The workflow of Anthropic's MCP.
Figure 10: The workflow of Google's A2A.
@@ III-C Communication Protocols for IoA
Traditionally, most agent frameworks only consider agents operating within their own ecosystems [7, 52, 41, 53, 37, 54, 47]. Such isolated architecture hinders interoperability, restricts the integration of third-party agents, and fails to reflect real-world scenarios where agents are distributed across heterogeneous devices and contexts. To overcome these fragmentation challenges, recent efforts focus on standardized agent communication protocols to support message exchange, task negotiation, context sharing, identity management, and integration with external tools, forming the backbone of scalable and interoperable IoA ecosystems. Representative agent communication protocols include the model context protocol (MCP) [13], agent-to-agent (A2A) [9], agent network protocol (ANP) [46], AGNTCY [111], and Agora [11], as summarized in Table IX.
III-C1 Anthropic's MCP
@@
III-C2 Google's A2A
Unlike the MCP, which focuses on connecting individual LLMs with external data sources and tools, Google's A2A protocol serves as a networking layer to facilitate seamless interoperability among heterogeneous AI agents. It facilitates direct communication and knowledge sharing for task collaboration. A2A incorporates enterprise-grade authentication and authorization mechanisms compliant with OpenAPI standards by default [9], such as OAuth 2.0 and OpenID Connect.
A2A facilitates communication between a client agent and a remote agent [9], as shown in Fig. 10. The client agent formulates and transmits tasks, while the remote agent responds by executing the tasks or providing relevant information. Agent collaboration begins when the client agent discovers a remote agent by retrieving its agent card, i.e., a standardized JSON file typically hosted at a well-known URL. This agent card contains essential metadata, including agent's capabilities, skills, endpoint address, and authentication requirements, enabling structured agents discovery across diverse systems. Then, the client agent initiates a task by sending a request containing a unique task ID via universally unique identifier (UUID).The remote agent handles the task according to its declared capabilities. For asynchronous tasks, intermediate server-sent events (SSE) are streamed as the task progresses, providing status updates and partial response. In contrast, synchronous tasks return the final result directly in the response. If additional input is needed, the client sends follow-up messages using the same task ID until the task reaches a terminal state (e.g., completed or failed).
@@ III-C3 Other Protocols
  • Mixture-of-Agents (MoA). MoA [114] enhances agent coordination emulating a ``collaborative LLM'' through a layered architecture, where each layer contains multiple LLM-based agents with the same model architecture that refine the previous layer's outputs as auxiliary context.
Summary and Lessons Learned: In IoA systems, standardized communication protocols are vital to enable interoperability among heterogeneous agents, laying the foundation for scalable and context-aware collaboration. Notably, MCP provides agents with structured and standardized access to external tools and resources through unified interfaces, while A2A emphasizes P2P interactions and dynamic agent discovery. In parallel, ANP focuses on decentralized networking; AGNTCY adopts a hybrid architecture to promote standardized inter-agent collaboration; Agora explores agent integration with a balance of efficiency, versatility, and portability; and MoA targets on task-oriented layered agent orchestration. Despite differences in architecture and functionality, these protocols share a common goal: to build thriving, extensible, and trustworthy IoA ecosystems. Table IX provides a comparative summary of their core goals, architectural designs, authentication mechanisms, agent discovery strategies, endorsing entities, advantages, and limitations.
@@ III-D Consensus & Conflict Resolution in IoA
2) How to Align Diverse Agent Reasoning? To overcome the inherent limitations of LLM agents, such as diverse reasoning paths and hallucination, reasoning alignment techniques extend beyond simple majority voting [63] to achieve more reliable consensus. Semantic interoperability remains a key challenge for the IoA, particularly as agents become increasingly heterogeneous in model architecture, training data, and tool usage. Such diversity often leads to issues such as sycophancy (agents reinforcing each other’s responses without critical reasoning), prompt ambiguity, and difficulty in reaching robust consensus.
@@
  • Dynamic Prompt Optimization: To improve alignment among LLM-based agents, CONSENSAGENT [116] introduces a dynamic prompt optimization mechanism to refine prompts based on agent interactions, which mitigates sycophancy and improves consensus efficiency in multi-agent LLM debates.
  • Collective Reasoning: Beyond prompt engineering, another promising direction is to utilize the collective power of multiple LLM agents to enhance semantic alignment, enabling mutual verification and augmentation. Such collective process enhances inference depth and robustness, particularly for complex, multi-hop reasoning tasks. For instance, He et al. [117] demonstrate that aggregating reasoning from diverse models can improve factual accuracy and approximate correctness, even without ground truth.
    a) Round-table discussion. In [78], the RECONCILE scheme leverages round-table discussions among diverse LLM agents, incorporating confidence-weighted voting and convincingness to enhance collaborative reasoning and consensus-building.
    b) Multi-agent deliberation. To further improve trustworthy negotiation, Yang et al. [118] present a training-free, multi-agent deliberation protocol where agents collectively calibrate their confidence estimates through argumentation, peer feedback, and majority voting to calibrate agent confidence and improve reliability in collective decision-making.
    c) MoA. To address the challenge of aligning agent outputs to specific task requirements at inference time, Chakraborty et al. [119] propose an MoA controlled decoding strategy, enabling inference-time alignment by dynamically selecting the most suitable agent for each token, thus facilitating fine-grained semantic interoperability without retraining.
  • Ontology Matching: For more structured semantic alignment tasks, such as ontology matching, Qiang et al. [120] propose an agent-powered LLM framework named Agent-OM that uses Siamese LLM agents to retrieve, match, and validate entity correspondences across heterogeneous ontologies, employing CoT reasoning and RAG to achieve semantic alignment between heterogeneous knowledge sources.
3) How to Scale Consensus in Large-Scale IoA? To mitigate scalability issues in consensus formation, hierarchical consensus mechanisms could be employed. Agents are partitioned into local clusters based on network topology or task similarity, with each cluster forms local consensus independently. Then, cluster representatives propagate aggregated results to higher-level coordinators, enabling efficient, scalable, and coherent system-wide decision-making. For instance, Feng et al. [121] propose hierarchical consensus framework that forms global consensus from agents' local observations via contrastive learning. By incorporating multi-layer consensus and an adaptive attention mechanism, the framework enhances decentralized cooperation and improves scalability.
@@ III-E1 Pricing Models
  • Contribution-aware Pricing rewards agents based on their marginal impact on task outcomes. A common approach employs the Shapley value to quantify each agent's marginal contribution. Specifically, agents with greater influence over critical task components receive proportionally larger shares of the collective reward, enabling fair profit distribution in multi-agent settings. For instance, Heuillet et al. [61] apply Monte Carlo sampling to estimate Shapley values in a MARL setting to infer and reward cooperative behavior in complex environments.
@@ III-E2 Economic Mechanism Design
TABLE X: Comparison of Incentive Mechanism Designs in IoA Systems
@@
a) Incentive Mechanism Design. Typically, agents are rewarded through monetary rewards, e.g., tradable tokens redeemable for computing resources, based on task performance [125]. Freni et al. [129] present a comprehensive model for agent-driven token economy design which integrates token-based incentives and governance tokens to align agent interests, support decentralized decision-making, and promote ecosystem sustainability. Prior works on incentives in MAS provides valuable lessons for IoA, including auction theory [126], contract theory [127], game theory [122, 123], and reputation-based mechanisms [128].
@@
  • Auction Theory. In auction mechanisms, agents bid for tasks based on resource availability and task valuation, enabling fair competition and efficient allocation. For instance, Yang et al. [130] introduce the concept of an agent-centric economy and propose Agent Exchange (AEX), i.e., a real-time auction platform where agents act as autonomous economic actors, bidding and negotiating for tasks and resources. The AEX framework leverages market-based protocols and dynamic value attribution (e.g., Shapley value) to ensure fair and efficient agent coordination. You et al. [126] further design a stochastic auction mechanism for efficient and privacy-preserving resource allocation among multiple agents.
@@
  • Contract Theory. By specifying performance-based contract items for agents with diverse types, contract-theoretic approaches provide a structured approach to mitigate information asymmetry and incentivize agents to align their actions with the objective of the principal (e.g., task coordinator agent). For instance, Ye et al. [127] employ contract theory and generative diffusion models to optimize multi-dimensional resource allocation under asymmetric information, offering insights for incentive design in IoA systems. Besides, Ivanov et al. [131] combine RL with principal-agent theory, illustrating that contract-based incentives can align agent behavior with system-level objectives in sequential and multi-agent environments.
@@
  • Game Theory. Game-theoretic approaches in IoA are divided into non-cooperative and cooperative models games. ① Non-cooperative games model agents as rational and selfish entities whose strategic interactions converge to a Nash equilibrium, at which no agent can unilaterally improve its payoff by deviating. For instance, Geng et al. [122] formulate a non-cooperative game among distributed quadrotor UAVs and derive a generalized Nash equilibrium of the game under communication and physical constraints. ② Cooperative games encourage agents to independently form coalitions, with coalitional benefits distributed fairly among its members via mechanisms such as Shapley values. For instance, Wang et al. [123] model social cluster formation as a cooperative game, where agents collaboratively optimize group strategies to enhance overall utility while ensuring equitable coalition benefit allocation. From a joint behavioral and game-theoretic perspective, Lalmohammed [132] develops a welfare modeling framework for human-AI economic ecosystems, incorporating trust evolution, risk aversion, and collaboration synergies. Using agent-based modeling, the study shows that dynamic trust-building and equitable incentive distribution are pivotal for maximizing collective welfare in mixed human-AI systems.
@@
  • Reputation mechanisms track agents' historical performance, prioritizing high-reputation agents for critical tasks and discouraging malicious behavior. For instance, Xu et al. [128] propose a hierarchical reputation framework that selects reliable agents based on reputation scores for secure collaboration in dynamic and uncertain environments. To further enable transparent and automated incentive mechanisms, Karim et al. [133] explore the integration of AI agents with blockchain, showing how smart contracts and decentralized ledgers facilitate agent-to-agent payments, reputation, and governance.
@@
  • Game-theoretic Penalties. Leveraging principles from game theory, several incentive-compatible penalty schemes have developed to align individual agent behavior with system-wide objectives. For instance, Xu et al. [134] design game-theoretical reward and penalty policies based on agents' reported beliefs and past performance, thereby improving task quality while reducing latency.
@@
Summary and Lessons Learned: Economic models in IoA align individual agent incentives with collective objectives to sustain long-term collaboration. In IoA pricing models, capability-based pricing charges agents according to resource usage (e.g., token usage), whereas contribution-aware pricing allocates rewards agents proportionally to their marginal contributions. However, accurately estimating contributions in large-scale and dynamic IoA environments is often computationally prohibitive. Beyond pricing, mechanism design integrates incentives and penalties to foster cooperation. Incentive mechanisms grounded in auction, contract, and game theories promote active participation, while reputation mechanisms support long-term collaboration. Table X summarizes representative incentive works on agents along with their strengths and limitations. Conversely, penalty mechanisms deter misbehaving agents by imposing costs (e.g., token forfeiture or reputation loss). A key challenge is developing interoperable and fair economic frameworks across heterogeneous IoA platforms with varying protocols and value models.